OpenStudy (anonymous):

10q - 20 is equivalent to _____. 1)2(q - 10) 2)10(q - 20) 3)10(q - 2)

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

you confused me

OpenStudy (anonymous):

help me please

OpenStudy (ryanwhite499):

What do you think

OpenStudy (anonymous):

i think the answer is number 1

OpenStudy (anonymous):

2(q-10)

OpenStudy (ryanwhite499):

Hmmmm... hold on

OpenStudy (anonymous):

i see now it is B 10(q-20)

OpenStudy (ryanwhite499):

Yes ...

OpenStudy (ryanwhite499):

Oh yes that is correct 10q-20=0 10q=20 q=20/10=2

OpenStudy (anonymous):

10(q-20) If you distribute the 10, will you get the original?

OpenStudy (anonymous):

what about this math problem: Simplify and determine the coefficient of (negative 2/3x)(3y)(-2x)

OpenStudy (anonymous):

@Worior9131 for your original problem, did you check the distribution to make sure it worked?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

really? cuz 10*20 = 200

OpenStudy (anonymous):

the other person said i was correct

OpenStudy (anonymous):

the other person was solving an equation. Your problem isn't an equation. It's asking you to factor/reverse distribute

OpenStudy (anonymous):

it is asking for the equivalent not what is = to 20

OpenStudy (anonymous):

equivalent means equal, so it has to equal 10q + 20 when you distribute

OpenStudy (ryanwhite499):

@peachpi she is talking about this problem. What are you talking about?

OpenStudy (ryanwhite499):

Or he

OpenStudy (anonymous):

sorry 10q-20

OpenStudy (anonymous):

i am a he (Acmed voice)

OpenStudy (anonymous):

\[10q-20 \neq 10(q-20)\]

OpenStudy (ryanwhite499):

Thats not on of his options @peachpi

OpenStudy (ryanwhite499):

He answered that correctly. He said it was B

OpenStudy (ryanwhite499):

Its correct

OpenStudy (anonymous):

you guys kill me stressing out about the options. LEARN THE MATH, and then you'll be able to see the correct option, which isn't B

OpenStudy (anonymous):